Please help me solve the following problem.
Interest rate. A credit manager rates each applicant for a car loan on a scale of 1 through 5 and than determines the interest rate from the accompanying table. Find the equation of the line in slope intercept form that goes through these points.
Credit Rating Interest Rate%
1 24
2 20
3 16
4 12
5 8
Thank you so much for your help

Answer by stanbon(75887)

1 24
2 20
3 16
4 12
5 8
The line will have the form y=mx+b.
Take any two of the paired numbers; solve for m, then for b, as follows:
If we choose (2,20) and (5,8) you get:
m=[(20-8)/(2-5)]=12/-3=-4
Substitute a value for x,y, and m into y=mx+b to find b, as follows:
20=(-4)(2)+b
b=28
So the equation you want is y=-4x+28
Cheers,
Stan H.